# On a generalisation of Cameron’s base size conjecture

### v2

@@ 4. Fixed point ratios for classical groups
Using [12, Tables 8.12-8.14], as well as Table 3, we can determine the possibilities of $H$. For most of those cases we use entirely similar methods as the ones we use to prove Propositions 4.9, 4.10, and 4.12, so we only discuss the cases where we use different methods here, to avoid unnecessary repetition. In particular, we discuss the cases where $H$ is of type $^{2}B_{2}(q)$, $O_{2}^{\epsilon}(q)\wr S_{2}$, or $O^{-}_{2}(q^{2}).2$ for $q$ even, ${\rm L}_{2}(q)$ for $q$ odd.
If $H$ is of type $^{2}B_{2}(q)$, then we obtain the required bounds by inspecting the proof of [18, 4.4], so it remains to discuss the other four cases.
Suppose that $H$ is of type $O_{2}^{\epsilon}(q)\wr S_{2}$. Note that if $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{6}$, then the trivial bound $|x^{G}\cap H|<|H|$ is sufficient. Now suppose that $|x^{G_{0}}|\leqslant\frac{1}{2}\left(\frac{q}{q+1}\right)q^{6}$. Lemma 3.18 implies that this only occurs when $x$ is a $b_{1}$ involution. In this case we know that $x\in G_{0}$, and so $x^{G}\cap H=x^{G}\cap H_{0}$, where $H_{0}=H\cap G_{0}$. We can infer from [12, Table 8.14] that $H_{0}<K<G_{0}$, where $K={\rm Sp}_{2}(q)\wr S_{2}$, and so $|x^{G}\cap H|=|x^{G}\cap H_{0}|\leqslant|x^{G}\cap K|$. Now [16, 2.9] gives $|x^{G}\cap K|\leqslant 2q^{2}$ and we get $|x^{G_{0}}|=q^{4}-1$, and one can check that those bounds are sufficient for $q>32$.
If $q\leqslant 32$, then we can construct $H_{0}$ using Magma as the normaliser of an appropriate Sylow subgroup. We only give the details for the case where $H$ is of type $\Omega^{-}_{2}(q)\wr S_{2}$ when $q=32$, as the other cases are similar. In this case we can obtain $H_{0}$ as the normaliser of a Sylow $3$-subgroup. Note that $3$ divides $q+1$, and in particular, the $3$-factor of $|G_{0}|$ is the same as the $3$-factor of $(q+1)^{2}$, so a Sylow $3$-subgroup $P$ of $G_{0}$ is contained in some maximal torus $T\cong C_{q+1}\times C_{q+1}$. In particular, $P\cong C_{3}\times C_{3}$ is a characteristic subgroup of $T$, so $N_{G}(T)\leqslant N_{G}(P)$. One then checks using Magma that $|N_{G}(P)|=|H_{0}|$, so $N_{G}(P)=H_{0}$, and we can then check that the bound $|x^{G}\cap H|<i_{2}(H_{0})$ is sufficient.
The case where $H$ is of type $O^{-}_{2}(q^{2}).2$ is very similar. If $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{6}$, then the trivial bound $|x^{G}\cap H|<|H|$ is sufficient, so it remains to prove the claim for $b_{1}$-type involutions. Again we note that $x\in G_{0}$, and so $x^{G}\cap H=x^{G}\cap H_{0}$, and this time $H_{0}=H\cap G_{0}$ embeds inside a maximal subgroup $K$ of type ${\rm Sp}_{2}(q^{2})$, so $|x^{G}\cap H|\leqslant|x^{G}\cap K|$. However, $K$ does not contain any $b_{1}$-type involutions, and therefore $|x^{G}\cap H|=|x^{G}\cap K|=0$, and the claim trivially holds.
@@
$|x^{G_{0}}|\geqslant\frac{|{\rm Sp}_{4}(q)|}{|{\rm GL}_{1}(q)|^{2}}=q^{4}(q+1)^{2}(q^{2}+1).$
The case where $i=2$ is similar, so we leave the details to the reader.
We finally consider involutions. Let $K$ denote the algebraic closure of $\mathbb{F}_{q}$ and let $\bar{G}={\rm PSp}_{4}(K)$ and $\bar{H}={\rm PSL}_{2}(K)$. Now let $\bar{W}$ be the natural module for ${\rm SL}_{2}(K)$ and $\bar{\rho}:{\rm SL}_{2}(K)\rightarrow{\rm Sp}_{4}(K)$ denote the representation afforded by the $K{\rm SL}_{2}(q)$-module $S^{3}(\bar{W})$. Now note that $\bar{H}$ has a unique class of involutions, $y^{\bar{H}}$, and $y$ lifts in ${\rm SL}_{2}(K)$ to a diagonal matrix $\hat{y}=[\lambda,\lambda^{-1}]$ for some $\lambda\in K^{\times}$ of order $4$. Then
@@
The methods we use here are similar to the ones in the proofs of Propositions 4.11 and 4.13, so we leave the details for the reader. ∎
@@
This follows by combining Propositions 4.3, 4.4, 4.5, 4.6, 4.7, 4.8, 4.9, 4.10, 4.11, 4.12, 4.13, 4.14, and 4.15. ∎
@@ 5. Eta function bounds
Our aim here is to derive a similar bound to the one used in Proposition 5.1. To do this, we must bound the number of $G_{0}$-conjugacy classes of prime order elements of each type in $G$, as well as the size of each $G_{0}$-conjugacy class, and then group them accordingly. We record the relevant bounds for each type of prime order element in Table 11, but for the sake of brevity, we only discuss a few cases in detail. The cases not discussed are very similar.
@@
We can estimate how many classes correspond to each Jordan decomposition by counting the number of labelled partitions corresponding to that form and consulting [22, 3.5.14] to determine which $G_{I}$-classes split in $G_{0}$. The $G_{0}$-class bounds in Table 11 are obtained via [15, 3.21].
@@
Firstly, we know that every $N\in\mathbb{N}$ has at most $\log(N+2)$ distinct prime divisors, and so $|\pi|\leqslant\log(\log_{p}q+2)$ if $q\neq p$ and $|\pi|=0$ if $p=q$. Moreover, [44, 4.9.1] gives $|\mathcal{F}_{r}|\leqslant(6,q-\epsilon)(r-1)\leqslant 6(\log_{p}q+1)$. Similarly, it follows straight from [22, 3.2.15] that the $G_{I}$-class of graph-field automorphisms splits into $(6,q^{1/2}-1)$ $G_{0}$-classes, and so $|\mathcal{F}_{gf}|\leqslant 6$.
We now consider unipotent elements. Note that $(8)$ in [15] tells us that there are at most $2^{s}$ unipotent ${\rm Inndiag}(G_{0})$-classes with $\nu(x)=s$ for $s\in\{1,\ldots,,5\}$ and each class can split to at most $6$ classes in $G_{0}$. Therefore, $k_{s,u}<6\cdot 2^{s}$. Moreover, we can infer from [15, 3.22] and [22, 3.2.7] that
@@
Similarly, for semisimple elements [15, 3.40] tells us that there are at most $q^{s+1}$ prime order elements of nu-value $s$ of order $r$, so now it suffices to bound the possibilities for $r$ to bound $k_{s,s}$. Note that $r$ must divide $\prod_{i=1}^{3}(q^{i}-\epsilon)(q^{2i}-1)<q^{22}$, and since every integer $N>6$ has at most $\log N$ divisors, we conclude that $k_{s,s}<22\log q\cdot q^{s+1}$.
@@
[…]
@@
Now let $x_{1},\ldots,x_{t}$ be a set of representatives of the prime order conjugacy classes of $G$, and without loss of generality assume that $x_{t-1}$ and $x_{t}$ are the representatives of the unipotent $G$-classes with Jordan form $J_{2}^{2},J_{1}^{3}]$ and $J_{3}^{2},J_{1}]$ respectively. Propositions 4.7 and 4.8 tell us that ${\rm fpr}(x_{j},G/H_{1})<|x_{j}^{G}|^{-0.4}$ for all $j\in[t]$, whilst ${\rm fpr}(x_{j},G/H_{k})<|x_{j}|^{-0.268}$ for $k\neq 1$ and $j\in[t-2]$. Finally, for $j\in\{t-1,t-2\}$ ${\rm fpr}(x_{j},G/H_{k})$ can be recovered from [17, Tables 2.4 and 2.5]. Combining all of this we get
@@
Moreover, noting that $\alpha,\beta\rightarrow 0$ as $q\rightarrow\infty$ and combining this with Proposition 3.13, we find that $\mathbb{P}(G,\tau)\rightarrow 1$ as $|G|\rightarrow\infty$. Finally, if $\sigma$ is a $6$-tuple of $G$ all of whose entries are of type $G_{2}(q)$, then [25, Theorem 3] implies that $\mathbb{P}(G,\sigma)\rightarrow 1$ as $|G|\rightarrow\infty$, and so $\mathbb{P}(G,6)\rightarrow 1$ as $|G|\rightarrow\infty$, as required.
@@
Let $\tau=(H_{1},\ldots,H_{4})$ be a non-standard tuple. If $G_{0}={\rm PSL}^{\epsilon}_{5}(q)$ or ${\rm PSL}^{\epsilon}_{4}(q)$, then in view of Propositions 4.9, 4.10, and 4.11, we may write
@@
where $\mathcal{F}$ denotes the set of prime order field and graph-field automorphisms in $G$ and $\mathcal{G}$ the set of involutory graph automorphisms in $G$. We will derive bounds for $\alpha_{i}$ and $\beta_{i}$ for $i\in[3]$ for each possibility for $G_{0}$ separately.
First suppose that $G_{0}={\rm PSL}^{\epsilon}_{5}(q)$. If $q\leqslant 4$, then the regularity of $\tau$ can be verified using Magma, so we may assume that $q\geqslant 5$. In this case, we can infer from [41, Table 3] that the total number of conjugacy classes in ${\rm PGL}_{5}^{\epsilon}(q)$ is bounded above by $q^{4}+8q^{3}$. We now note that the only $G$-class of unipotent elements that potentially splits in $G_{0}$ is the one corresponding to the Jordan form $J_{5}]$, and [22, 3.2.7, 3.3.10] imply that it splits into at most 55 $G_{0}$-classes. Therefore, we may take $\alpha_{1}\leqslant q^{4}+8q^{3}+4$. Moreover, [44, 4.9.1] implies that there are at most
$\left(5,\frac{q-\epsilon}{q^{1/r}-\epsilon}\right)(r-1)|\pi(\log_{p}(q))|\leqslant 5\log_{p}q\cdot\log(\log_{p}q+2)$
$G_{0}$-classes of field automorphisms and similarly at most $5$ distinct $G_{0}$-classes of involutory graph-field automorphisms by [22, 3.2.15]. We also get at most 55 $G_{0}$-classes of involutory graph automorphisms by [22, 3.2.14, 3.3.17], so we have
@@
Next assume that $G_{0}={\rm PSL}^{\epsilon}_{4}(q)$. If $q\leqslant 9$, then the regularity of $\tau$ can again be verified using Magma, so we may assume that $q\geqslant 11$ for the remainder. Again, from [41, Table 3] we see that there are at most $\alpha_{1}\leqslant q^{3}+8q^{2}$ $G_{I}$-classes of unipotent and semisimple elements, and the only classes that possibly split in $G_{0}$ are the class of unipotent elements corresponding to the Jordan decompositions $J_{4}]$ and $J_{2}^{2}]$, so $\alpha_{1}\leqslant q^{3}+8q^{2}+4$. Moreover, [44, 4.9.1] again tells us that the contribution to $\alpha_{2}$ from field automorphisms is at most $4\log_{p}q\cdot\log(\log_{p}q+2)$, and [22, 3.2.15] tells us that there are at most 44 $G_{0}$-classes of involutory graph-field automorphisms. Finally, we can infer from [22, 3.2.14, 3.3.17] that each $G_{I}$-class of graph automorphisms splits into at most 22 $G_{0}$-classes, and so
@@
Then Propositions 4.12 and 4.13 give us
@@
so it suffices to find sufficient bounds for $\alpha_{i},\beta_{i}$ for $i\in[4]$. We can again infer from [41, Table 3] that there are at most $q^{2}+8q$ distinct prime order classes in $G_{I}$, none of which split in $G_{0}$, so $\alpha_{1}\leqslant q^{2}+8q$. Moreover, the same arguments we used when $G_{0}={\rm PSL}_{5}(q)$ gives $\alpha_{2}\leqslant 3\log_{p}q\cdot\log(\log_{p}q+2)$, $\alpha_{3}\leqslant 3$ and $\alpha_{4}\leqslant 6$. Next, as usual we obtain
@@
Finally, suppose that $G_{0}={\rm PSp}_{4}(q)$. We verify the claim using Magma if $q\leqslant 11$ so we may assume that $q\geqslant 13$ for the remainder. Let $\alpha_{i},\beta_{i}$ be defined as in the case $G_{0}={\rm PSL}_{3}^{\epsilon}(q)$ for $i\in[4]$. Then Propositions 4.14 and 4.15 imply that (6) still applies, so we will again derive sufficient bounds for $\alpha_{i},\beta_{i}$ for $i\in[4]$.
By inspecting [41, Table 3] we deduce that there are at most $q^{2}+12q$ distinct $G_{I}$-classes of unipotent and semisimple elements and each unipotent class can split into at most 44 $G_{0}$-classes, so we get $\alpha_{1}\leqslant q^{2}+12q+9$. Moreover, Lemma 3.18 implies that $\beta_{1}>\frac{1}{4}\left(\frac{q}{q+1}\right)q^{4}$. Next, using [22, 3.4.15] we find
@@
and once again via Lemma 3.18, we get $\beta_{2}>\frac{1}{4}\left(\frac{q}{q+1}\right)q^{20/3}$. Similarly, we deduce that there are at most 22 $G_{0}$-classes of involutory field automorphisms by inspecting [22, 3.4.15], and [22, 3.4.16] tells us that there is an additional $G_{0}$ class of involutory graph-field automorphisms when $p=2$, so $\alpha_{4}\leqslant 3$. Finally, we get $\beta_{4}>\frac{1}{4}q^{5}$ from Lemma 3.18, and trivially, $\alpha_{3}=\beta_{3}=0$, since $G$ contains no graph automorphisms. One can check that those bounds are sufficient for $q\geqslant 13$.
@@
7. Preliminaries on exceptional groups

### v3

@@ 4. Fixed point ratios for classical groups
Using [12, Tables 8.12-8.14], as well as Table 3, we can determine the possibilities of $H$. For most of those cases we use entirely similar methods as the ones we use to prove Propositions 4.10, 4.11, and 4.13, so we only discuss the cases where we use different methods here, to avoid unnecessary repetition. In particular, we discuss the cases where $H$ is of type $^{2}B_{2}(q)$, $O_{2}^{\epsilon}(q)\wr S_{2}$, or $O^{-}_{2}(q^{2}).2$ for $q$ even, and ${\rm L}_{2}(q)$ for $q$ odd.
If $H$ is of type $^{2}B_{2}(q)$, then we obtain the required bounds by inspecting the proof of [18, 4.4], so it remains to handle the other four cases.
Suppose that $H$ is of type $O_{2}^{\epsilon}(q)\wr S_{2}$. Note that if $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{6}$, then the trivial bound $|x^{G}\cap H|<|H|$ is sufficient. Now suppose that $|x^{G_{0}}|\leqslant\frac{1}{2}\left(\frac{q}{q+1}\right)q^{6}$. Lemma 3.18 implies that this occurs when $x$ is an $a_{2}$ or $b_{1}$ involution. In these cases we note that $x\in G_{0}$, and so $x^{G}\cap H=x^{G}\cap H_{0}$, where $H_{0}=H\cap G_{0}$. We now note that $H_{0}=C_{q-\epsilon}^{2}{:}{\rm D}_{8}$, and in particular, $i_{2}(H_{0})=(q+\epsilon)(q+3)$. We therefore compute
$|x^{G_{I}}\cap H|\leqslant i_{2}(H_{0})\leqslant(q+1)(q+3)\leqslant(q^{4}-1)^{11/20}=|x^{G_{0}}|^{11/20}$
when $q>32$. If $q\leqslant 32$, then we can construct $H_{0}$ using Magma as the normaliser of an appropriate Sylow subgroup. We only give the details for the case where $H$ is of type $\Omega^{-}_{2}(q)\wr S_{2}$ when $q=32$, as the other cases are similar. In this case we can obtain $H_{0}$ as the normaliser of a Sylow $3$-subgroup. Note that $3$ divides $q+1$, and in particular, the largest power of $3$ dividing $|G_{0}|$ is the same as the largest power of $3$ dividing $(q+1)^{2}$, so a Sylow $3$-subgroup $P$ of $G_{0}$ is contained in some maximal torus $T\cong C_{q+1}\times C_{q+1}$. In particular, $P\cong C_{3}\times C_{3}$ is a characteristic subgroup of $T$, so $N_{G}(T)\leqslant N_{G}(P)$. One then checks using Magma that $|N_{G}(P)|=|H_{0}|$, so $N_{G}(P)=H_{0}$, and we can verify that the bound $|x^{G}\cap H|<i_{2}(H_{0})$ is sufficient.
The case where $H$ is of type $O^{-}_{2}(q^{2}).2$ is very similar. If $q\leqslant 32$, then we again use Magma, so now assume that $q\geqslant 64$. If $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{6}$, then the trivial bound $|x^{G}\cap H|<|H|$ is sufficient, so it remains to prove the claim for $a_{2}$ and $b_{1}$-type involutions. Again we note that $x\in G_{0}$, and so $x^{G}\cap H=x^{G}\cap H_{0}$. We now note that $H_{0}=C_{q^{2}+1}:4$, and in particular $i_{2}(H_{0})=q^{2}+1$. We then have
$|x^{G_{I}}\cap H|\leqslant i_{2}(H_{0})=q^{2}+1\leqslant(q^{4}-1)^{11/20}=|x^{G_{0}}|^{11/20}$
as required.
@@
$|x^{G_{0}}|\geqslant\frac{|{\rm Sp}_{4}(q)|}{|{\rm GL}_{1}(q)|^{2}}=q^{4}(q+1)^{2}(q^{2}+1),$
which is sufficient. The case where $i=2$ is similar, so we leave the details to the reader.
We finally consider involutions. Let $K$ denote the algebraic closure of $\mathbb{F}_{q}$ and let $\bar{G}={\rm PSp}_{4}(K)$ and $\bar{H}={\rm PSL}_{2}(K)$. Now let $\bar{W}$ be the natural module for ${\rm SL}_{2}(K)$ and $\bar{\rho}:{\rm SL}_{2}(K)\rightarrow{\rm Sp}_{4}(K)$ denote the representation afforded by the $K{\rm SL}_{2}(q)$-module $S^{3}(\bar{W})$. Now note that $\bar{H}$ has a unique class of involutions $y^{\bar{H}}$, and $y$ lifts in ${\rm SL}_{2}(K)$ to a diagonal matrix $\hat{y}=[\lambda,\lambda^{-1}]$ for some $\lambda\in K^{\times}$ of order $4$. Then
@@
The methods we use here are similar to the ones in the proofs of Propositions 4.12 and 4.14, so we leave the details for the reader. ∎
@@
This follows by combining Propositions 4.3, 4.4, 4.5, 4.6, 4.8, 4.9, 4.10, 4.11, 4.12, 4.13, 4.14, 4.15, and 4.16. ∎
@@ 5. Eta function bounds
Our aim here is to derive a similar bound to the one used in Proposition 5.1. To do this, we must bound the number of $G_{0}$-conjugacy classes of prime order elements of each type in $G$, as well as the size of each $G_{0}$-conjugacy class. We record the relevant bounds for each type of prime order element in Table 12, but for the sake of brevity, we only discuss a few cases in detail. The cases not discussed are very similar.
@@
We can estimate how many classes correspond to each Jordan decomposition by counting the number of signed partitions corresponding to that form and consulting [22, 3.5.14] to determine which $G_{I}$-classes split in $G_{0}$. The $G_{0}$-class bounds in Table 12 are obtained via [15, 3.21].
@@
Firstly, we know that every $N\in\mathbb{N}$ has at most $\log(N+2)$ distinct prime divisors, and so $|\pi|\leqslant\log(\log_{p}q+2)$ if $p\neq q$ and $|\pi|=0$ if $p=q$. Moreover, [45, 4.9.1] gives $|\mathcal{F}_{r}|\leqslant(6,q-\epsilon)(r-1)\leqslant 6(\log_{p}q+1)$. Similarly, it follows straight from [22, 3.2.15] that the $G_{I}$-class of graph-field automorphisms splits into $(6,q^{1/2}-1)$ $G_{0}$-classes, and so $|\mathcal{F}_{gf}|\leqslant 6$.
We now consider unipotent elements. Note that $(8)$ in [15] tells us that there are at most $2^{s}$ unipotent ${\rm Inndiag}(G_{0})$-classes with $\nu(x)=s$ for $s\in\{1,\ldots,5\}$ and each class can split to at most $6$ classes in $G_{0}$. Therefore, $k_{s,u}<6\cdot 2^{s}$. Moreover, we can infer from [15, 3.22] and [22, 3.2.7] that
@@
Similarly, for semisimple elements [15, 3.40] tells us that there are at most $q^{s+1}$ prime order elements with $\nu(x)=s$ of order $r$, so now it suffices to bound the possibilities for $r$ to bound $k_{s,s}$. Note that $r$ must divide $\prod_{i=1}^{3}(q^{i}-\epsilon)(q^{2i}-1)<q^{22}$, and since every integer $N>6$ has at most $\log N$ divisors, we conclude that $k_{s,s}<22\log q\cdot q^{s+1}$.
@@
Finally assume that $i=1$, so $r=7$. Since $7$ does not divide $6$, we find that $\hat{x}=[\lambda I_{2},\lambda^{\alpha}I_{2},\lambda^{\beta}I_{2}]$ for some non-trivial $7$-th root of unity $\lambda$ and $1<\alpha,\beta\leqslant 6$, so there are $\binom{7}{3}=35$ possibilities for $x$. Therefore $k_{4,s}=24+1+35=90\ll 22\log 8\cdot 8^{5}$. The argument in all remaining cases is similar, so we omit the details. ∎
@@
Now let $x_{1},\ldots,x_{t}$ be a set of representatives of the prime order conjugacy classes of $G$, and without loss of generality assume that $x_{t-1}$ and $x_{t}$ are the representatives of the unipotent $G$-classes with Jordan form $J_{2}^{2},J_{1}^{3}]$ and $J_{3}^{2},J_{1}]$ respectively. Propositions 4.8 and 4.9 tell us that ${\rm fpr}(x_{j},G/H_{1})<|x_{j}^{G}|^{-0.4}$ for all $j\in[t]$, whilst ${\rm fpr}(x_{j},G/H_{k})<|x_{j}|^{-0.268}$ for $k\neq 1$ and $j\in[t-2]$. Finally, for $j\in\{t-1,t-2\}$ ${\rm fpr}(x_{j},G/H_{k})$ can be recovered from [17, Tables 2.4 and 2.5]. Combining all of this we get
@@
Moreover, noting that $\alpha,\beta\rightarrow 0$ as $q\rightarrow\infty$ and combining this with Proposition 3.13, we find that $\mathbb{P}(G,\tau)\rightarrow 1$ as $|G|\rightarrow\infty$. Finally, if $\sigma$ is a $4$-tuple of $G$ all of whose entries are of type $G_{2}(q)$, then it follows immediately from [23, Theorem 5] that $\mathbb{P}(G,\sigma)\rightarrow 1$ as $|G|\rightarrow\infty$, and so $\mathbb{P}(G,4)\rightarrow 1$ as $|G|\rightarrow\infty$, as required.
@@
Let $\tau=(H_{1},\ldots,H_{4})$ be a non-standard tuple. If $G_{0}={\rm PSL}^{\epsilon}_{5}(q)$ or ${\rm PSL}^{\epsilon}_{4}(q)$, then in view of Propositions 4.10, 4.11, and 4.12, we may write
@@
where $\mathcal{F}$ denotes the set of prime order field and graph-field automorphisms in $G$ and $\mathcal{G}$ the set of involutory graph automorphisms in $G$. We will derive bounds for each $\alpha_{i}$ and $\beta_{i}$.
First suppose that $G_{0}={\rm PSL}^{\epsilon}_{5}(q)$. If $q\leqslant 4$, then the regularity of $\tau$ can be verified using Magma, so we may assume that $q\geqslant 5$. In this case, we can infer from [42, Table 3] that the total number of conjugacy classes in ${\rm PGL}_{5}^{\epsilon}(q)$ is bounded above by $q^{4}+8q^{3}$. We now note that the only $G$-class of unipotent elements that potentially splits in $G_{0}$ is the one corresponding to the Jordan form $J_{5}]$, and [22, 3.2.7, 3.3.10] imply that it splits into at most 55 $G_{0}$-classes. Therefore, we may take $\alpha_{1}\leqslant q^{4}+8q^{3}+4$. Moreover, [45, 4.9.1] implies that there are at most
$\left(5,\frac{q-\epsilon}{q^{1/r}-\epsilon}\right)(r-1)\leqslant 5\log_{p}q$
$G_{0}$-classes of field automorphisms of prime order $r$, and since $|\pi(\log_{p}(q))|\leqslant\log(\log_{p}q+2)$, there are at most
$5\log_{p}q\cdot\log(\log_{p}q+2)$
distinct $G_{0}$-classes of prime order field automorphisms in $G$. Similarly, there are at most $5$ distinct $G_{0}$-classes of involutory graph-field automorphisms by [22, 3.2.15]. We also get at most $5$ distinct $G_{0}$-classes of involutory graph automorphisms by [22, 3.2.14, 3.3.17], so we have
@@
Next assume that $G_{0}={\rm PSL}^{\epsilon}_{4}(q)$. If $q\leqslant 9$, then the regularity of $\tau$ can again be verified using Magma, so we may assume that $q\geqslant 11$ for the remainder. Again, from [42, Table 3] we see that there are at most $\alpha_{1}\leqslant q^{3}+8q^{2}$ distinct $G_{I}$-classes of unipotent and semisimple elements, and the only classes that possibly split in $G_{0}$ are the class of unipotent elements corresponding to the Jordan decompositions $J_{4}]$ and $J_{2}^{2}]$, so $\alpha_{1}\leqslant q^{3}+8q^{2}+4$. Moreover, [45, 4.9.1] again tells us that the contribution to $\alpha_{2}$ from field automorphisms is at most $4\log_{p}q\cdot\log(\log_{p}q+2)$, and [22, 3.2.15] tells us that there are at most $4$ distinct $G_{0}$-classes of involutory graph-field automorphisms. Finally, we can infer from [22, 3.2.14, 3.3.17] that each $G_{I}$-class of graph automorphisms splits into at most 22 $G_{0}$-classes, and so
@@
Then Propositions 4.13 and 4.14 give us
@@
so it suffices to find sufficient bounds for $\alpha_{i},\beta_{i}$ for $i\in[4]$. We can again infer from [42, Table 3] that there are at most $q^{2}+8q$ distinct prime order classes in $G_{I}$, none of which split in $G_{0}$, so $\alpha_{1}\leqslant q^{2}+8q$. Moreover, the same arguments we used when $G_{0}={\rm PSL}_{5}(q)$ gives $\alpha_{2}\leqslant 3\log_{p}q\cdot\log(\log_{p}q+2)$, $\alpha_{3}\leqslant 3$ and $\alpha_{4}\leqslant 6$. Next, as usual we obtain
@@
Finally, suppose that $G_{0}={\rm PSp}_{4}(q)$. We verify the claim using Magma if $q\leqslant 11$ so we may assume that $q\geqslant 13$ for the remainder. Let $\alpha_{i},\beta_{i}$ be defined as in the case $G_{0}={\rm PSL}_{3}^{\epsilon}(q)$ for $i\in[4]$. Then Propositions 4.15 and 4.16 imply that (8) still applies, so we will again derive sufficient bounds for $\alpha_{i},\beta_{i}$ for $i\in[4]$.
By inspecting [42, Table 3] we deduce that there are at most $q^{2}+12q$ distinct $G_{I}$-classes of unipotent and semisimple elements and each unipotent class can split into at most $4$ distinct $G_{0}$-classes, so we get $\alpha_{1}\leqslant q^{2}+12q+9$. Moreover, Lemma 3.18 implies that $\beta_{1}>\frac{1}{4}\left(\frac{q}{q+1}\right)q^{4}$. Next, using [22, 3.4.15] we find
@@
and once again via Lemma 3.18, we get $\beta_{2}>\frac{1}{4}\left(\frac{q}{q+1}\right)q^{20/3}$. Similarly, we deduce that there are at most $2$ distinct $G_{0}$-classes of involutory field automorphisms by inspecting [22, 3.4.15], and [22, 3.4.16] tells us that there is an additional $G_{0}$-class of involutory graph-field automorphisms when $p=2$, so $\alpha_{4}\leqslant 3$. Finally, we get $\beta_{4}>\frac{1}{4}q^{5}$ from Lemma 3.18, and trivially, $\alpha_{3}=\beta_{3}=0$, since $G$ contains no graph automorphisms. One can check that those bounds are sufficient for $q\geqslant 13$.
@@
7. Exceptional group preliminaries